The common stock of Kyocera currently sells for $88.50 and its current dividend is $1.10. Determine...

  1. The common stock of Kyocera currently sells for $88.50 and its current dividend is $1.10. Determine the growth rate for Kyocera assuming that an investor's required rate of return is 14% and that earnings and dividends are expected to grow at a constant rate.

    A.

    13.9%    

    B.

    12.3%    

    C.

    13.8%

    D.

    12.6%
Answer #1

Required return=(D1/Current price)+Growth rate

0.14=(1.1*(1+Growth rate))/88.5+Growth rate

0.14=(1.1+1.1*Growth rate)/88.5+Growth rate

0.14=0.0124293785+0.0124293785*Growth rate+Growth rate

Growth rate=(0.14-0.0124293785)/(0.0124293785+1)

=12.6%(Approx).